\begin{thebibliography}{43}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Afridi et~al.(2021)Afridi, Alam, Khan, Khan, and
  Lee}]{afridi2021multimodal}
Tariq~Habib Afridi, Aftab Alam, Muhammad~Numan Khan, Jawad Khan, and Young~Koo
  Lee. 2021.
\newblock A multimodal memes classification: A survey and open research issues.
\newblock In \emph{5th International Conference on Smart City Applications, SCA
  2020}, pages 1451--1466. Springer Science and Business Media Deutschland
  GmbH.

\bibitem[{Alam et~al.(2021{\natexlab{a}})Alam, Alam, Hasan, Hasnat, Imran, Ofli
  et~al.}]{alam2021medic}
Firoj Alam, Tanvirul Alam, Md~Hasan, Abul Hasnat, Muhammad Imran, Ferda Ofli,
  et~al. 2021{\natexlab{a}}.
\newblock {MEDIC:} a multi-task learning dataset for disaster image
  classification.
\newblock \emph{arXiv:2108.12828}.

\bibitem[{Alam et~al.(2021{\natexlab{b}})Alam, Alam, Imran, and
  Ofli}]{alam2021social}
Firoj Alam, Tanvirul Alam, Muhammad Imran, and Ferda Ofli. 2021{\natexlab{b}}.
\newblock Robust training of social media image classification models for rapid
  disaster response.
\newblock \emph{arXiv:2104.04184}.

\bibitem[{Alam et~al.(2021{\natexlab{c}})Alam, Cresci, Chakraborty, Silvestri,
  Dimitrov, Martino, Shaar, Firooz, and Nakov}]{alam2021survey}
Firoj Alam, Stefano Cresci, Tanmoy Chakraborty, Fabrizio Silvestri, Dimiter
  Dimitrov, Giovanni Da~San Martino, Shaden Shaar, Hamed Firooz, and Preslav
  Nakov. 2021{\natexlab{c}}.
\newblock A survey on multimodal disinformation detection.
\newblock \emph{arXiv:2103.12541}.

\bibitem[{Ben-younes et~al.(2019)Ben-younes, Cadene, Thome, and
  Cord}]{BlockFusion2019}
Hedi Ben-younes, Remi Cadene, Nicolas Thome, and Matthieu Cord. 2019.
\newblock \href
  {https://ojs.aaai.org/index.php/AAAI/article/download/4818/4691}
  {Block:bilinear superdiagonal fusion for visual question answeringand visual
  relationship detection}.
\newblock In \emph{Proceedings of the 33st Conference on Artificial
  Intelligence (AAAI)}, volume~15.

\bibitem[{Bondielli and Marcelloni(2019)}]{bondielli2019survey}
Alessandro Bondielli and Francesco Marcelloni. 2019.
\newblock A survey on fake news and rumour detection techniques.
\newblock \emph{Information Sciences}, 497:38--55.

\bibitem[{Brooke(2019)}]{brooke-2019-condescending}
Sian Brooke. 2019.
\newblock \href {https://doi.org/10.18653/v1/W19-3519} {{``Condescending, Rude,
  Assholes''}: Framing gender and hostility on {S}tack {O}verflow}.
\newblock In \emph{Proceedings of the Third Workshop on Abusive Language
  Online}, pages 172--180, Florence, Italy. Association for Computational
  Linguistics.

\bibitem[{Chandra et~al.(2021)Chandra, Pailla, Bhatia, Sanchawala, Gupta,
  Shrivastava, and Kumaraguru}]{chandra2021subverting}
Mohit Chandra, Dheeraj Pailla, Himanshu Bhatia, Aadilmehdi Sanchawala, Manish
  Gupta, Manish Shrivastava, and Ponnurangam Kumaraguru. 2021.
\newblock {``Subverting the Jewtocracy''}: Online antisemitism detection using
  multimodal deep learning.
\newblock \emph{arXiv:2104.05947}.

\bibitem[{Chen et~al.(2019)Chen, Li, Yu, El~Kholy, Ahmed, Gan, Cheng, and
  Liu}]{chen2019uniter}
Yen-Chun Chen, Linjie Li, Licheng Yu, Ahmed El~Kholy, Faisal Ahmed, Zhe Gan,
  Yu~Cheng, and Jingjing Liu. 2019.
\newblock {UNITER}: Learning universal image-text representations.

\bibitem[{Da~San~Martino et~al.(2020)Da~San~Martino, Cresci,
  Barr{\'{o}}n{-}Cede{\~{n}}o, Yu, Pietro, and Nakov}]{da2020survey}
Giovanni Da~San~Martino, Stefano Cresci, Alberto Barr{\'{o}}n{-}Cede{\~{n}}o,
  Seunghak Yu, Roberto~Di Pietro, and Preslav Nakov. 2020.
\newblock \href {https://doi.org/10.24963/ijcai.2020/672} {A survey on
  computational propaganda detection}.
\newblock In \emph{Proceedings of the Twenty-Ninth International Joint
  Conference on Artificial Intelligence, {IJCAI} 2020}, pages 4826--4832.
  ijcai.org.

\bibitem[{Da~San~Martino et~al.(2019)Da~San~Martino, Yu, Barr{\'o}n-Cede{\~n}o,
  Petrov, and Nakov}]{EMNLP19DaSanMartino}
Giovanni Da~San~Martino, Seunghak Yu, Alberto Barr{\'o}n-Cede{\~n}o, Rostislav
  Petrov, and Preslav Nakov. 2019.
\newblock \href {https://doi.org/10.18653/v1/D19-1565} {Fine-grained analysis
  of propaganda in news article}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 5636--5646, Hong Kong,
  China. Association for Computational Linguistics.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and Toutanova}]{text_bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1423} {{BERT}: Pre-training of
  deep bidirectional transformers for language understanding}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Dimitrov et~al.(2021{\natexlab{a}})Dimitrov, Bin~Ali, Shaar, Alam,
  Silvestri, Firooz, Nakov, and Da~San~Martino}]{dimitrov2021detecting}
Dimitar Dimitrov, Bishr Bin~Ali, Shaden Shaar, Firoj Alam, Fabrizio Silvestri,
  Hamed Firooz, Preslav Nakov, and Giovanni Da~San~Martino. 2021{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.516} {Detecting
  propaganda techniques in memes}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing}, ACL-IJCNLP~'21, pages 6603--6617, Online.
  Association for Computational Linguistics.

\bibitem[{Dimitrov et~al.(2021{\natexlab{b}})Dimitrov, Bin~Ali, Shaar, Alam,
  Silvestri, Firooz, Nakov, and Da~San~Martino}]{SemEval2021-6-Dimitrov}
Dimitar Dimitrov, Bishr Bin~Ali, Shaden Shaar, Firoj Alam, Fabrizio Silvestri,
  Hamed Firooz, Preslav Nakov, and Giovanni Da~San~Martino. 2021{\natexlab{b}}.
\newblock {Task 6 at SemEval-2021}: Detection of persuasion techniques in texts
  and images.
\newblock In \emph{Proceedings of the 15th International Workshop on Semantic
  Evaluation}, SemEval~'21, Bangkok, Thailand. Association for Computational
  Linguistics.

\bibitem[{Fersini et~al.(2019)Fersini, Gasparini, and
  Corchs}]{fersini2019detecting}
Elisabetta Fersini, Francesca Gasparini, and Silvia Corchs. 2019.
\newblock Detecting sexist meme on the web: A study on textual and visual cues.
\newblock In \emph{2019 8th International Conference on Affective Computing and
  Intelligent Interaction Workshops and Demos (ACIIW)}, pages 226--231. IEEE.

\bibitem[{Fortuna and Nunes(2018)}]{fortuna2018survey}
Paula Fortuna and S{\'e}rgio Nunes. 2018.
\newblock A survey on automatic detection of hate speech in text.
\newblock \emph{ACM Computing Surveys (CSUR)}, 51(4):1--30.

\bibitem[{Gan et~al.(2020)Gan, Chen, Li, Zhu, Cheng, and Liu}]{gan2020large}
Zhe Gan, Yen-Chun Chen, Linjie Li, Chen Zhu, Yu~Cheng, and Jingjing Liu. 2020.
\newblock Large-scale adversarial training for vision-and-language
  representation learning.
\newblock \emph{Advances in Neural Information Processing Systems},
  33:6616--6628.

\bibitem[{Gasparini et~al.(2021)Gasparini, Rizzi, Saibene, and
  Fersini}]{france_mys2021}
Francesca Gasparini, Giulia Rizzi, Aurora Saibene, and Elisabetta Fersini.
  2021.
\newblock \href {http://arxiv.org/abs/2106.08409} {Benchmark dataset of memes
  with text transcriptions for automatic detection of multi-modal misogynistic
  content}.
\newblock \emph{CoRR}, abs/2106.08409.

\bibitem[{Haidar et~al.(2016)Haidar, Chamoun, and Yamout}]{7920246}
Batoul Haidar, Maroun Chamoun, and Fadi Yamout. 2016.
\newblock \href {https://doi.org/10.1109/EMS.2016.037} {Cyberbullying
  detection: A survey on multilingual techniques}.
\newblock In \emph{2016 European Modelling Symposium (EMS)}, pages 165--171.

\bibitem[{He et~al.(2016)He, Zhang, Ren, and Sun}]{he2016deep}
Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016.
\newblock Deep residual learning for image recognition.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 770--778.

\bibitem[{Hegde et~al.(2021)Hegde, Hande, Priyadharshini, Thavareesan, and
  Chakravarthi}]{Siddhanth2021}
Siddhanth~U Hegde, Adeep Hande, Ruba Priyadharshini, Sajeetha Thavareesan, and
  Bharathi~Raja Chakravarthi. 2021.
\newblock Uvce-iiitt@ dravidianlangtech-eacl2021: Tamil troll meme
  classification: You need to pay more attention.
\newblock In \emph{Proceedings of the First Workshop on Speech and Language
  Technologies for Dravidian Languages}, pages 180--186.

\bibitem[{Husain and Uzuner(2021)}]{husain2021survey}
Fatemah Husain and Ozlem Uzuner. 2021.
\newblock A survey of offensive language detection for the arabic language.
\newblock \emph{ACM Transactions on Asian and Low-Resource Language Information
  Processing (TALLIP)}, 20(1):1--44.

\bibitem[{Kiela et~al.(2021)Kiela, Firooz, Mohan, Goswami, Singh, Fitzpatrick,
  Bull, Lipstein, Nelli, Zhu et~al.}]{kiela2021hateful}
Douwe Kiela, Hamed Firooz, Aravind Mohan, Vedanuj Goswami, Amanpreet Singh,
  Casey~A Fitzpatrick, Peter Bull, Greg Lipstein, Tony Nelli, Ron Zhu, et~al.
  2021.
\newblock The hateful memes challenge: competition report.
\newblock In \emph{NeurIPS 2020 Competition and Demonstration Track}, pages
  344--360. PMLR.

\bibitem[{Kiela et~al.(2020)Kiela, Firooz, Mohan, Goswami, Singh, Ringshia, and
  Testuggine}]{kiela2020hateful}
Douwe Kiela, Hamed Firooz, Aravind Mohan, Vedanuj Goswami, Amanpreet Singh,
  Pratik Ringshia, and Davide Testuggine. 2020.
\newblock \href
  {https://proceedings.neurips.cc/paper/2020/hash/1b84c4cee2b8b3d823b30e2d604b1878-Abstract.html}
  {The hateful memes challenge: Detecting hate speech in multimodal memes}.
\newblock In \emph{Advances in Neural Information Processing Systems}, Online.

\bibitem[{Kingma and Ba(2015)}]{kingma2014adam}
Diederik~P. Kingma and Jimmy Ba. 2015.
\newblock Adam: {A} method for stochastic optimization.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Kolesnikov et~al.(2021)Kolesnikov, Dosovitskiy, Weissenborn, Heigold,
  Uszkoreit, Beyer, Minderer, Dehghani, Houlsby, Gelly, Unterthiner, and
  Zhai}]{Dosovitskiy2021}
Alexander Kolesnikov, Alexey Dosovitskiy, Dirk Weissenborn, Georg Heigold,
  Jakob Uszkoreit, Lucas Beyer, Matthias Minderer, Mostafa Dehghani, Neil
  Houlsby, Sylvain Gelly, Thomas Unterthiner, and Xiaohua Zhai. 2021.
\newblock An image is worth 16x16 words: Transformers for image recognition at
  scale.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Kumar et~al.(2018)Kumar, Ojha, Malmasi, and
  Zampieri}]{kumar2018benchmarking}
Ritesh Kumar, Atul~Kr Ojha, Shervin Malmasi, and Marcos Zampieri. 2018.
\newblock Benchmarking aggression identification in social media.
\newblock In \emph{Proceedings of the First Workshop on Trolling, Aggression
  and Cyberbullying}, pages 1--11.

\bibitem[{Li et~al.(2019)Li, Yatskar, Yin, Hsieh, and Chang}]{li2019visualbert}
Liunian~Harold Li, Mark Yatskar, Da~Yin, Cho-Jui Hsieh, and Kai-Wei Chang.
  2019.
\newblock {VisualBERT}: A simple and performant baseline for vision and
  language.
\newblock \emph{arXiv:1908.03557}.

\bibitem[{Mubarak et~al.(2017)Mubarak, Darwish, and Magdy}]{mubarak2017abusive}
Hamdy Mubarak, Kareem Darwish, and Walid Magdy. 2017.
\newblock Abusive language detection on arabic social media.
\newblock In \emph{Proceedings of the first workshop on abusive language
  online}, pages 52--56.

\bibitem[{Nakov et~al.(2021)Nakov, Giovanni, Elsayed,
  Barr{\'{o}}n{-}Cede{\~{n}}o, M\'{i}guez, Shaar, Alam, Haouari, Hasanain,
  Mansour, Hamdan, Ali, Babulkov, Nikolov, Shahi, Struß, Mandl, Kutlu, and
  Kartal}]{clef-checkthat:2021:LNCS}
Preslav Nakov, Da~San~Martino Giovanni, Tamer Elsayed, Alberto
  Barr{\'{o}}n{-}Cede{\~{n}}o, Rub\'{e}n M\'{i}guez, Shaden Shaar, Firoj Alam,
  Fatima Haouari, Maram Hasanain, Watheq Mansour, Bayan Hamdan, Zien~Sheikh
  Ali, Nikolay Babulkov, Alex Nikolov, Gautam~Kishore Shahi, Julia~Maria
  Struß, Thomas Mandl, Mucahid Kutlu, and Yavuz~Selim Kartal. 2021.
\newblock Overview of the {CLEF}-2021 {CheckThat}! lab on detecting
  check-worthy claims, previously fact-checked claims, and fake news.
\newblock LNCS (12880). Springer.

\bibitem[{Ofli et~al.(2020)Ofli, Alam, and Imran}]{multimodalbaseline2020}
Ferda Ofli, Firoj Alam, and Muhammad Imran. 2020.
\newblock Analysis of social media data using multimodal deep learning for
  disaster response.
\newblock In \emph{Proceedings of the Information Systems for Crisis Response
  and Management}.

\bibitem[{Platt(1998)}]{platt1998sequential33}
John Platt. 1998.
\newblock Sequential minimal optimization: A fast algorithm for training
  support vector machines.

\bibitem[{Pramanick et~al.(2021)Pramanick, Sharma, Dimitrov, Akhtar, Nakov, and
  Chakraborty}]{pramanick-etal-2021-momenta-multimodal}
Shraman Pramanick, Shivam Sharma, Dimitar Dimitrov, Md.~Shad Akhtar, Preslav
  Nakov, and Tanmoy Chakraborty. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.findings-emnlp.379}
  {{MOMENTA}: A multimodal framework for detecting harmful memes and their
  targets}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EMNLP 2021}, pages 4439--4455, Punta Cana, Dominican Republic. Association
  for Computational Linguistics.

\bibitem[{Simonyan and Zisserman(2015)}]{simonyan2014very}
Karen Simonyan and Andrew Zisserman. 2015.
\newblock \href {http://arxiv.org/abs/1409.1556} {Very deep convolutional
  networks for large-scale image recognition}.
\newblock In \emph{3rd International Conference on Learning Representations},
  an Diego, CA, USA.

\bibitem[{Smith(2007)}]{smith2007overview}
Ray Smith. 2007.
\newblock An overview of the tesseract ocr engine.
\newblock In \emph{Ninth international conference on document analysis and
  recognition}, volume~2, pages 629--633.

\bibitem[{Su et~al.(2019)Su, Zhu, Cao, Li, Lu, Wei, and Dai}]{su2019vl}
Weijie Su, Xizhou Zhu, Yue Cao, Bin Li, Lewei Lu, Furu Wei, and Jifeng Dai.
  2019.
\newblock {VL-BERT}: Pre-training of generic visual-linguistic representations.
\newblock \emph{arXiv:1908.08530}.

\bibitem[{Suryawanshi and Chakravarthi(2021)}]{dravidiantrollmeme-eacl}
Shardul Suryawanshi and Bharathi~Raja Chakravarthi. 2021.
\newblock Findings of the shared task on {T}roll {M}eme {C}lassification in
  {T}amil.
\newblock In \emph{Proceedings of the First Workshop on Speech and Language
  Technologies for Dravidian Languages}. Association for Computational
  Linguistics.

\bibitem[{Suryawanshi et~al.(2020{\natexlab{a}})Suryawanshi, Chakravarthi,
  Arcan, and Buitelaar}]{suryawanshi-etal-2020-multimodal}
Shardul Suryawanshi, Bharathi~Raja Chakravarthi, Mihael Arcan, and Paul
  Buitelaar. 2020{\natexlab{a}}.
\newblock \href {https://aclanthology.org/2020.trac-1.6} {Multimodal meme
  dataset ({M}ulti{OFF}) for identifying offensive content in image and text}.
\newblock In \emph{Proceedings of the Second Workshop on Trolling, Aggression
  and Cyberbullying}, pages 32--41, Marseille, France. European Language
  Resources Association (ELRA).

\bibitem[{Suryawanshi et~al.(2022)Suryawanshi, Chakravarthi, Arcan, Levy,
  Buitaleer, Kumaresan, Ponnusamy, and Hande}]{trollmeme-eacl}
Shardul Suryawanshi, Bharathi~Raja Chakravarthi, Mihael Arcan, Susan Levy, Paul
  Buitaleer, Prasanna~Kumar Kumaresan, Rahul Ponnusamy, and Adeep Hande. 2022.
\newblock Findings of the second shared task on {T}roll {M}eme {C}lassification
  in {T}amil.
\newblock In \emph{Proceedings of the Second Workshop on Speech and Language
  Technologies for Dravidian Languages}. Association for Computational
  Linguistics.

\bibitem[{Suryawanshi et~al.(2020{\natexlab{b}})Suryawanshi, Chakravarthi,
  Verma, Arcan, McCrae, and Buitelaar}]{suryawanshi-etal-2020-dataset}
Shardul Suryawanshi, Bharathi~Raja Chakravarthi, Pranav Verma, Mihael Arcan,
  John~Philip McCrae, and Paul Buitelaar. 2020{\natexlab{b}}.
\newblock \href {https://www.aclweb.org/anthology/2020.wildre-1.2} {A dataset
  for troll classification of {T}amil{M}emes}.
\newblock In \emph{Proceedings of the WILDRE5{--} 5th Workshop on Indian
  Language Data: Resources and Evaluation}, pages 7--13, Marseille, France.
  European Language Resources Association (ELRA).

\bibitem[{Tan and Le(2019)}]{tan2019efficientnet}
Mingxing Tan and Quoc~V Le. 2019.
\newblock {EfficientNet:} rethinking model scaling for convolutional neural
  networks.
\newblock \emph{arXiv:1905.11946}.

\bibitem[{Van~Hee et~al.(2015)Van~Hee, Lefever, Verhoeven, Mennes, Desmet,
  De~Pauw, Daelemans, and Hoste}]{van-hee-etal-2015-detection}
Cynthia Van~Hee, Els Lefever, Ben Verhoeven, Julie Mennes, Bart Desmet, Guy
  De~Pauw, Walter Daelemans, and Veronique Hoste. 2015.
\newblock \href {https://aclanthology.org/R15-1086} {Detection and fine-grained
  classification of cyberbullying events}.
\newblock In \emph{Proceedings of the International Conference Recent Advances
  in Natural Language Processing}, pages 672--680, Hissar, Bulgaria. INCOMA
  Ltd. Shoumen, BULGARIA.

\bibitem[{Zia et~al.(2021)Zia, Castro, and Tyson}]{zia-etal-2021-racist}
Haris~Bin Zia, Ignacio Castro, and Gareth Tyson. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.woah-1.23} {Racist or sexist
  meme? classifying memes beyond hateful}.
\newblock In \emph{Proceedings of the 5th Workshop on Online Abuse and Harms
  (WOAH 2021)}, pages 215--219, Online. Association for Computational
  Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[]{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}
\usepackage{natbib}
\usepackage{microtype}
\usepackage{booktabs}
\usepackage{subcaption}
\usepackage{url}
\usepackage{comment}
\usepackage{todonotes}
\usepackage{color,soul}
\usepackage{graphicx}
\usepackage[normalem]{ulem}
\usepackage{algorithm}
\usepackage{algorithmic}
\usepackage{amsmath,amssymb}
\usepackage{longtable}
\usepackage{lscape}
\usepackage{enumitem}
\usepackage[normalem]{ulem}
\useunder{\uline}{\ul}{}

\title{TeamX@DravidianLangTech-ACL2022: A Comparative Analysis for Troll-Based Meme Classification}

\author{Rabindra Nath Nandi$^1$, Firoj Alam$^2$, Preslav Nakov$^2$\\
  $^1$BJIT Limited, Dhaka, Bangladesh \\
  $^2$Qatar Computing Research Institute, HBKU, Doha, Qatar \\
  \texttt{rabindra.nath@bjitgroup.com},
  \texttt{\{falam,pnakov\}@hbku.edu.qa}
  \\}

\begin{document}
\maketitle
\begin{abstract}
The spread of fake news, propaganda, misinformation, disinformation, and harmful content online raised concerns among social media platforms, government agencies, policymakers, and society as a whole. This is because such harmful or abusive content leads to several consequences to people such as physical, emotional, relational, and financial. Among different harmful content \textit{trolling-based} online content is one of them, where the idea is to post a message that is provocative, offensive, or menacing with an intent to mislead the audience. The content can be textual, visual, a combination of both, or a meme. In this study, we provide a comparative analysis of troll-based memes classification using the textual, visual, and multimodal content. We report several interesting findings in terms of code-mixed text, multimodal setting, and combining an additional dataset, which shows improvements over the majority baseline. 
\end{abstract}
\section{Introduction}
\label{sec:intro}
Social media have become one of the main communication channels for the propagation of information through textual, visual, or audio-visual content. While the content shared on social media creates a positive impact, however, there are also content that spread harm and hostility \citep{brooke-2019-condescending}, 
including 
abusive language \citep{mubarak2017abusive}, propaganda \cite{da2020survey,EMNLP19DaSanMartino}
cyberbullying \citep{van-hee-etal-2015-detection}, 
cyber-aggression \citep{kumar2018benchmarking}, and other kinds of harmful content \citep{pramanick-etal-2021-momenta-multimodal}.
The propagation of such content most often is done by an automated tool, troll, or coordinated groups, which target specific users, communities (e.g., minority groups), individuals, and companies. To detect such content there has been effort to develop automatic tools (see most recent surveys on 
disinformation \citep{alam2021survey}, rumours \citep{bondielli2019survey}, propaganda \citep{da2020survey}, and multimodal memes  \citep{afridi2021multimodal}, hate speech \citep{fortuna2018survey}, 
cyberbullying \citep{7920246}, and offensive content \citep{husain2021survey}. In addition, shared tasks has also been organized in the past years addressing factuality, fake news and harmful content~\cite{clef-checkthat:2021:LNCS,kiela2020hateful}. 

Among other social media content, recently, the uses of \textit{internet memes} became popular and they are often shared for the purpose of humor or fun with no bad intentions. However, memes are also created and shared with harmful intentions. This include attack on people based on the characteristics such as ethnicity, race, sex, gender identity, disability, disease, nationality, and immigration status \cite{kiela2020hateful}.
There has been research effort to develop computational method to detect such memes, such as detecting hateful memes \cite{kiela2020hateful}, propaganda \cite{dimitrov2021detecting,SemEval2021-6-Dimitrov}, offensive \cite{suryawanshi-etal-2020-multimodal}, sexist meme \cite{fersini2019detecting} and troll based meme \cite{dravidiantrollmeme-eacl}. 

In this study, we focus on troll-based meme classification based on the dataset released in the shared task discussed in \cite{trollmeme-eacl}. While meme contains both textual and visual elements, hence, we investigate textual, visual content and their combination
using different pretraind transformer models.
In addition, we explored combining an external dataset, and use of code-mixed text (i.e., Tamil and English) extracted using OCR. Note that the text provided with the dataset is transcribed in Latin. While prior work focuses on the text provided with the dataset, here, we also follow a different strategy, directly using the text from the OCR without any cleaning. 

Our contributions include:
\begin{itemize} 
\item 
we investigate classical algorithm (e.g., SVM), pretraind transformer and deep CNN models for both text and images, respectively;
\item we combine an additional dataset and use code-mixed text, extracted using OCR and compare the performance;  
\item we also experiment with different pretrained multimodal models. 
\end{itemize}
    
\section{Related Work}
\label{sec:related_work}
Prior work on detecting harmful aspects of memes include categorizing hateful memes \cite{kiela2020hateful}, antisemitism \cite{chandra2021subverting} and propaganda detection techniques in memes \cite{dimitrov2021detecting}, harmful memes and their target \cite{pramanick-etal-2021-momenta-multimodal}, identifying protected category such as race, sex  that has been attacked~\cite{zia-etal-2021-racist}, and identifying offensive content \cite{suryawanshi-etal-2020-multimodal}. Among the studies most notable effort that streamlined the research work include shared tasks such as ``Hateful Memes Challenge'' \cite{kiela2020hateful}, detection of persuasion techniques \cite{SemEval2021-6-Dimitrov} and troll meme classification \cite{dravidiantrollmeme-eacl}. 

The work by \citet{chandra2021subverting} investigates antisemitism along with its types by addressing the tasks as binary and multi-class classification using  pretrained transformers and CNN as modality-specific encoders along with various multimodal fusion strategies. \citet{dimitrov2021detecting} developed a dataset with 22 propaganda techniques and investigates the different state-of-the-art pretrained models and demonstrate that joint vision-language models perform best. \citet{pramanick-etal-2021-momenta-multimodal} address two tasks such as detecting harmful memes and identifying the social entities they target and propose a multimodal model, which utilizes local and global information. \citet{zia-etal-2021-racist} goes one step further than a binary classification of hateful memes -- more fine-grained categorization based on protected category (i.e., race, disability, religion, nationality, sex) and their attack types (i.e., contempt, mocking, inferiority, slurs, exclusion, dehumanizing, inciting violence) using the dataset released in the WOAH 2020 Shared Task.\footnote{\url{github.com/facebookresearch/fine_grained_hateful_memes}} 
\citet{fersini2019detecting} studied sexist meme detection and investigate textual cues with a late-fusion strategy, which suggest that fusion approach performs better. The same authors also developed a dataset of 800 misogynistic memes covering different manifestations of hatred against women (e.g., body shaming, stereotyping, objectification and violence), which are collected from different social media~\citep{france_mys2021}.

In the ``Hateful Memes Challenge'', the participants addressed the hateful meme classification task by fine-tuning the state-of-art multi-modal transformer models \citep{kiela2021hateful} and best system in the competition used different unimodal and multimodal pre-training models such as VisualBERT \citep{li2019visualbert} VL-BERT \citep{su2019vl}, UNITER \citep{chen2019uniter}, VILLA \citep{gan2020large} and 
ensembles \citep{kiela2021hateful}. 
The SemEval-2021 propaganda detection shared task \citep{SemEval2021-6-Dimitrov} was organized with a focus on fine-grained propaganda techniques in text and the entire meme, and from the participants' systems, they conclude that multimodal cues are important for automated propaganda detection. In the troll meme classification shared task \cite{dravidiantrollmeme-eacl}, the best system used ResNet152, BERT with multimodal attention, and the majority of the system used pretrained transformer models for text, CNN models for images, and early fusion approaches.

\section{Experiments}
\label{ssec:classification}
\subsection{Data}
\label{ssec:data}
We use the dataset provided in the troll-based Tamil meme classification shared task discussed in \cite{suryawanshi-etal-2020-dataset,trollmeme-eacl}. The dataset is comprised of meme and transcribed text in Latin, which are annotated and transcribed by native Tamil speakers. There is a total of 2,300, 667 memes for training and testing, respectively. For our experiments, we split the training set into training and development set with 80 and 20\%, respectively. The development set is used for fine-tuning the models.

In Table \ref{tab:data_dist}, we report the distribution of the dataset that we used for the experiments. 
\begin{table}[tbh!]
\centering
\begin{tabular}{lrrr}
\toprule
\textbf{Class } & \textbf{Train} & \textbf{Dev} & \textbf{Test} \\ \midrule
Troll & 1013  & 269 & 395 \\
Non-Troll & 827 & 191 & 272 \\
\bottomrule
\end{tabular}
\caption{Distribution of the Troll-based Tamil Meme dataset. We split original training set into training and development set.}
\label{tab:data_dist} 
\end{table}

\subsection{Settings}
\label{ssec:models}
For the classification, we run different unimodal 
experiments: {\em (i)} only text, {\em (ii)} only meme, and {\em (iii)} text and meme together. 
For each setting, we also run several baseline experiments. One such baseline is the \textit{majority class} baseline, which predicts the label based on the most frequent label in the training set. 
This has been most commonly used in shared tasks \cite{clef-checkthat:2021:LNCS}.
Furthermore, we run a few advanced experiments using an additional dataset and code-mixed text from OCR. To measure the performance of each model we used a weighted F$_1$ score to maintain shared task guideline.

\subsubsection{Text Modality}
For the baseline using text modality, we used bag-of-$n$-gram vectors weighted with logarithmic term frequencies (tf) multiplied with inverse document frequencies (idf) and train the model using Support Vector Machines (SVM)~\cite{platt1998sequential33}. Note that we extracted unigram, bigram, and tri-gram features. We used grid search to optimize the SVM hyper-parameters.

We then experiment using multilingual BERT (mBERT) model \cite{text_bert}.
We performed ten reruns for each experiment using different random seeds, then we picked the model that performed best on the development set. We used a batch size of 8, a learning rate of 2e-5, maximum sequence length 128, three epochs, and used the `categorical cross-entropy' as the loss function. 

\subsubsection{Image Modality}
Similar to the text modality, for the baseline experiment with image modality, we extract features from a pre-trained model, then train the model using SVM. We extracted features from the penultimate layer of the EfficientNet (b1) model \cite{tan2019efficientnet}, which was trained using ImageNet. For training the model using SVM we used the default parameters setting.

For the later experiments we used the transfer learning approach, fine-tuning the pre-trained deep CNN models (e.g., VGG16), which has been shown success for visual recognition tasks. 
We used the weights of the model pre-trained on ImageNet to initialize our model. We adapt the last layer (i.e., softmax layer) of the network for the binary classification task. 
We trained models using three popular neural network architectures such as VGG16~\cite{simonyan2014very}, ResNet101~\cite{he2016deep} and EfficientNet~\cite{tan2019efficientnet}, which showed state-of-art performance in similar tasks \cite{multimodalbaseline2020,alam2021medic,alam2021social}. For training, we used the Adam optimizer~\cite{kingma2014adam} with an initial learning rate of $10^{-5}$, which is decreased by a factor of 10 when accuracy on the dev set stops improving for 10 epochs. 
From our experiment, we observe that the model converse within 50-60 epochs. As the dataset size is small, fine-tuning the entire network did not yield better results, therefore, we freeze the network and fine-tune the penultimate layer.

\subsubsection{Multimodal: Text and Image}

For the multimodel experiments, we used Vision Transformer (ViT) \cite{Dosovitskiy2021} for image feature encoding and multilingual BERT (mBERT) for the textual representation. 
We used BLOCK fusion \cite{BlockFusion2019} to merge the features from two different modalities. 
BLOCK fusion is a multimodal fusion based on block-superdiagonal tensor decomposition. Previously, \citet{Dosovitskiy2021} showed a better performance using BLOCK fusion over several bilinear fusion techniques for Visual Question Answering (VQA) and Visual Relationship Detection (VRD) tasks. We conduct two major experiments by varying the textual data, {\em(i)} using the provided text, and {\em(ii)} using the code-mixed text.

\subsubsection{Additional Experiments} \mbox{} \\
\textbf{Code-mixed Tamil and English text:} 
The dataset comes with extracted text in a transcribed form in Latin. Given that current multilingual transformer models have not trained using such a latin form of text, therefore, to further understand the problem we extract text from memes using \textit{tesseract} \cite{smith2007overview}.\footnote{\url{https://pypi.org/project/pytesseract/}} 
We then train the same mBERT model using the extracted code-mixed Tamil and English text. Note that the extracted text contains noise that comes from the output of the OCR. 


\textbf{Additional data:} 
While we visually inspected the dataset we realized that textual and visual elements of memes have similarities with memes in hateful meme dataset \cite{kiela2020hateful}. We then mapped the labels from hateful memes dataset to troll labels, {\em(i)} \textit{hateful} to \textit{troll}, and {\em(ii)} \textit{not-hateful} to \textit{non-troll}. We combined all training and test set memes from the hateful memes dataset with the training set of the troll meme dataset. The dev set of the hateful meme dataset is combined with the dev set of the troll dataset. After combining the datasets we experiment with both text and image modalities using the same models.

\section{Results and Discussion}
\label{sec:discussion}

\begin{table}[]
\centering
\setlength{\tabcolsep}{2.0pt}
\scalebox{0.90}{
\begin{tabular}{lrrrr}
\toprule
\multicolumn{1}{c}{\textbf{Exp}} & \multicolumn{1}{c}{\textbf{Acc}} & \multicolumn{1}{c}{\textbf{P}} & \multicolumn{1}{c}{\textbf{R}} & \multicolumn{1}{c}{\textbf{F1}} \\ \midrule
Maj. & 59.2 & 35.1 & 59.2 & 44.1 \\ \midrule
\multicolumn{5}{c}{\textbf{Text modality}} \\ \midrule
Tf-Idf + SVM & 46.2 & 47.9 & 46.2 & \textbf{46.6} \\
mBERT & 52.6 & 51.0 & 52.6 & \textbf{51.4} \\
Add. data + mBERT & 56.5 & 52.8 & 56.5 & \textbf{51.8} \\
Code-mixed text + mBERT & 57.3 & 55.2 & 57.3 & \textbf{55.2} \\ \midrule
\multicolumn{5}{c}{\textbf{Image modality}} \\ \midrule
EffNet feat +SVM & 59.1 & 55.6 & 59.1 & \textbf{50.1} \\
VGG16 & 55.8 & 50.2 & 55.8 & \textbf{49.2} \\
ResNet101 & 60.4 & 58.6 & 60.4 & \textbf{53.8} \\
EffNet (b1) & 61.5 & 61.5 & 61.5 & \textbf{53.6} \\ 
\midrule
\multicolumn{5}{c}{\textbf{Multimodality}} \\ \midrule
ViT + mBERT & 57.9 & 55.1 & 57.9 & \textbf{54.2} \\
ViT + mBERT (code-mixed text) & 55.3 & 57.0 & 55.3 & \textbf{55.7} \\ \midrule
\multicolumn{5}{c}{\textbf{Image modality + Additional data}} \\ \midrule
VGG16 & 54.4 & 52.0 & 54.4 & \textbf{52.3} \\
ResNet101 & 60.7 & 58.9 & 60.7 & \textbf{56.2} \\
EffNet (b1) & 60.7 & 58.9 & 60.7 & \underline{\textbf{56.6}} \\ 
\bottomrule
\end{tabular}
}
\caption{Evaluation results on the test set. The results that improve over the majority class baseline are in \textbf{bold}, and the best one is \underline{underlined}. Maj.: Majority baseline, Add. data: Additional data.
}
\label{tab:results}
\end{table}

In Table \ref{tab:results}, we present the results for different modalities and settings. Overall, all results are better than majority baseline. In the unimodal experiments, the image-only models perform better than the text-only models. Our experiments on text-only models suggest that code-mixed data help in improving the performance using the same multiligual mBERT model compared to the transcribed latin text, which suggest that multilingual model is not able to capture the information in latin Tamil text. The additional data, which is in English only, slightly improved the performance for text-only experiment. 

For the image-only experiments, we obtain a comparative performance with ResNet101 and EfficientNet (EffNet (b1)). 


Our experiments on multimodel experiments provide better results compared to text and image modalities as shown in the Table \ref{tab:results}.
For the two multimodel experiments, we obtained weighted-F1 scores of 54.2 and 55.7 for text and for code-mixed text, respectively, which are better than the best weighted-F1 score from text modality (51.4) and from image modality (53.8).
Note that, the performance of our multimodel experiments is better than \citet{Siddhanth2021}, where a weighted F1 score of 0.47 has been reported for a similar task. That implies BLOCK fusion has advantages over late fusion for troll meme classification tasks. 

With additional data, results improved significantly for all image-only models. It confirms our observation that visual elements in memes share information across different datasets, and possibly different cultures too.

Our future plan is to apply the data augmentation technique for multimodality experiments and also we have a plan to explore the performance of popular multimodel algorithms.

During the shared task participation, we submitted one run, using only image modality, where our system was not showing promising performance. However, our subsequent experiments and analysis show several promising directions in terms of original code-mixed data and additional data from the other task.

\section{Conclusion and Future Work}
\label{sec:conclutions}
We present a comparative analysis of different modalities for the troll-based meme classification task. We show that the multilingual model capture more information for code-mixed text than its Latin counterpart. We present higher performance with multimodality compared to unimodal models. Our experiments also suggest that additional data from the other task helps in capturing visual information. In the future, we plan to further develop multimodal models that can capture information in code-mixed noisy conditions. 

\bibliography{bib/all_bib,bib/bibliography}
\bibliographystyle{acl_natbib}



\end{document}